OpenStudy (anonymous):

Plans for a bridge are drawn on a coordinate grid. One girder of the bridge lies on the line y = –9x – 9. A perpendicular brace passes through the point (–1, 7). Write an equation of the line that contains the brace.

OpenStudy (aroub):

y=mx+b , m=1/9 , (-1,7) 7=1/9(-1)+b 7=-1/9+b 64/9=b => y=1/9x+64/9